# Leveraging Vision-Language Models for Visual Grounding and Analysis of Automotive UI

### v1

@@ 3.4 Model Fine-Tuning
For efficient training of baseline models, we utilized LoRa [12]. LoRa offers benefits such as accelerated training, reduced training data needs, smaller storage footprint for model artifacts (adapters), lower VRAM requirements, and the potential for dynamic adapter swapping in deployment, eliminating the need to load the complete model.
To establish a suitable baseline model for our fine-tuning and experiments, we evaluated four existing VLMs: TinyClick (0.27B) [24], Molmo-7B-D-0924 [4], UGround-V1-7B (Qwen2-VL-based) [10], and InternVL2_5-8B [2]. TinyClick and UGround are fine-tuned versions of Florence-2-Base [34] and Qwen2-VL [30], respectively. These models are specifically designed to process UI images, interpret actions, and generate corresponding point coordinates. However, their broader capabilities beyond this visual grounding task are limited. They frequently output points even when the prompt requests different information. TinyClick’s relatively small size further restricts its potential as a robust foundation for our work. In contrast, Molmo and InternVL2.5 are general-purpose VLMs. Notably, they were also trained on datasets including GUI data from mobile and web applications. While InternVL2.5 can generate bounding boxes, Molmo is limited to point generation. When evaluated on our proposed AutomotiveUI-Bench-4k dataset, Molmo-7B-D-0924 achieved the best performance (Table 3) among the considered models and was therefore chosen as the foundation of ELAM.
We generated a dataset of 17,708 annotations for 6,230 images using the synthetic data pipeline described in Section 3.3. It comprises 5,952 Test Actions, 6,190 passed, and 5,566 failed evaluations with corresponding Expected Results. The images in our training datset were captured using 4k cameras, either from our vehicle fleet or with the generous consent of local car dealerships. This dataset was subsequently utilized for LoRa fine-tuning of the Molmo-7B-D-0924 checkpoint. The prompt templates used for training, evaluation, and inference are described in Table 1.
Training was conducted for 2 epochs with a global batch size of 128, achieved using a local batch size of 8 across two NVIDIA H100 80 GB PCIe GPUs. We employed a learning rate of 1e-4 and incorporated LoRa to enhance parameter efficiency, applying it to all linear layers within Molmo.
Table 1: Prompt and response templates for training and inference of ELAM as Python f-strings
Prompt Template - Test Action
Response Template - Test Action
Prompt Template - Expected Result
Response Template - Expected Result
@@ 4 Open Source Infotainment Validation Dataset: AutomotiveUI-Bench-4K
Given that multiple datasets [25, 33, 32, 6, 38, 1, 8] already exist in the fields of desktop, mobile, and web, we identify a domain gap in scientific research for benchmarking LAMs for automotive UI. To fill this gap, we introduce and release an infotainment UI dataset named AutomotiveUI-Bench-4K. This dataset comprises 998 images with 4,208 annotations from 15 brands such as Audi, BMW, Ford, Porsche, and Tesla among others, thus ensuring a diverse variety of modern vehicle UI, meticulously crafted by experienced professional software testers specializing in HMI evaluation. In addition to images from automotive OEMs, the dataset also incorporates screenshots of Apple CarPlay and Google Android Auto. Images from OEMs were captured in 4K resolution using cameras and sourced from either our vehicle fleet or through collaboration with local car dealerships. To ensure consistent image perspective and focus on the relevant user interface elements, the OEM images were rectified and cropped to the active display area. Android Auto and CarPlay screenshots were directly exported. The infotainment systems featured in this dataset are exclusively touchscreen-based. While some systems include auxiliary physical buttons, these were not considered in our analysis due to their primarily static functionality. Although the dataset primarily comprises systems from recent model years, the system representing the earliest model year is a 2018 Volkswagen Tiguan. Languages that were included are German and English. Instructions and evaluation requests are written in English. German text was either translated or directly quoted. Real-world test engineering documentation often exhibits a pattern of short sentences and occasional grammatical deviations. Therefore, to create a dataset that realistically represents this domain, it is crucial to preserve these stylistic features in the preprocessing steps. Table 2 shows further information about the data distribution. The actions and results are unrelated, since test actions can lead to different screens and we do not consider trajectories. There are two annotation classes in the dataset:
@@
    Test Action: Describes an interaction with a single UI element as an imperative sentence (”set A/C to max” in Fig. 3).
@@
    Expected Result: Represents a testable expectation, defined per image, focusing on the required state or behavior of a particular UI element within that image. This expectation is evaluated to determine a passed/failed outcome. (”Passenger’s climate zone is synced to driver” in Fig. 3)
@@ 5.1 Evaluating Performance with AutomotiveUI-Bench-4K
The performance of various baseline models and fine-tuned models are shown in Table 3. The numeric values represent accuracy in percent. $TA$ and $ER$ represent data samples for classes Test Action and Expected Results, respectively. Visual grounding scores are indicated by the index $vg$. Evaluations pertaining to predicted outcomes are denoted by the index $evl$. Accuracy was assessed by analyzing the generated point or centroid of the bounding box and verifying its containment within the annotated ground truth box. To assess language-specific performance, the image dataset was stratified into two distinct categories based on the language (English or German) represented in the images. When a generated evaluation could not be parsed from the model’s response, it was assigned the inverse of the ground truth value. All evaluation experiments for AutomotiveUI-Bench-4K were conducted using a NVIDIA RTX 4090 GPU with 24 GB VRAM.
The findings suggest that models within the 7B parameter range exhibit no significant performance degradation when processing images with predominantly German text. A significant challenge for most baseline models was the generation of consistently parseable output. For example, InternVL2.5 had good evaluation performance but rarely generated bounding boxes. TinyClick and UGround were only evaluated for their grounding performance as they do not have evaluative capabilities. To explore the possibility of faster inference through task distribution across models, we experimented with LoRa-fine-tuning TinyClick (LAM-270M), due to its size of only 0.27 billion parameters. This approach, however, did not produce performance on par with Molmo-7B as shown in Table 3, and we therefore abandoned further investigation in this direction.
Table 3: Evaluation on AutomotiveUI-Bench-4K
@@ 5.2 Evaluating Generalizability with ScreenSpot
To determine if our fine-tuning resulted in overfitting to the automotive UI domain, we applied ELAM-7B to the ScreenSpot dataset [3] comparing it to SeeClick [3], ShowUI [18], OS-Atlas-Base-7B [33], Molmo-7B-D-0924/Molmo-72B-0924 [4], and Qwen2-VL-based UGround-V1-7B [10] in Table 4. Our fine-tuning procedure did not break the generalizability of Molmo. The average score per category was improved by 5.2% compared to the baseline (Molmo-7B-D-0924) and by 2.2% compared to Molmo-72B.
Table 4: Evaluation results on ScreenSpot for selected models taken from [11]
@@ 5.3 Ablation Study
We performed ablation studies to evaluate the contributions of simple reasoning and rephrasing modules integrated into our synthetic data generation pipeline. Training runs were conducted using distinct preprocessing strategies, the configurations of which are summarized in Table 5. The prompts from Table 1 were used for experiments that incorporate reasoning, whereas the phrase “Think step by step, conclude” was replaced by “Determine” for runs that did not utilize reasoning. A LoRa-rank of 64 was chosen to fine-tune all linear layers for the ablation experiments. Our findings demonstrate that the inclusion of reasoning significantly enhances performance, as evidenced by improvements in both visual grounding and evaluation accuracy. This observation highlights the critical role of reasoning for robust generalization to real-world data in potential future deployments.
To optimize the LoRa configuration, we conducted a separate ablation study focusing on the rank parameter $r$, the results of which are presented in Table 6. The data was preprocessed utilizing reasoning and rephrasing. This analysis determined that a rank of 64 yielded optimal performance. In all LoRa-based experiments, the alpha parameter was set to $r$, and the LoRa-dropout rate was fixed at $0.05$.
@@
To investigate the relationship between semantic representation of user utterances and VLM performance, particularly within car application subdomains, we analyzed two distinct task types: Test Actions and Expected Results as outlined in Section 4. For both types, we generated text embeddings of the utterances using the nomic-embed-text-v1.5***nomic-embed-text-v1.5 is an improved variant of the nomic-embed-text-v1[23] model. It utilizes Matryoshka Representation Learning for flexible dimensionality reduction with minimal loss while supporting up to 8,192 tokens. The model is optimized for search, clustering, and classification in production. model and visualized their structure using t-SNE dimensionality reduction [27]. Unsupervised k-Means clustering ($k=8$) was applied to the high-dimensional embeddings to identify potential thematic groups, with cluster labels generated post-hoc via LLM (GPT-4o) interpretation of sampled utterances. Crucially, we enhanced the t-SNE projections with a heatmap that visualizes the local failure rate for each designated grounding for Test Actions and both verification and grounding for Expected Results. This failure rate was calculated per grid cell based on evaluation results, with red indicating high failure density and white indicating low failure density. This approach allows us to visualize how VLM performance, for both the base Molmo-7B and ELAM-7B models, varies across the fixed semantic landscape defined by the text embeddings.
@@ 5.4.1 Test Action Utterances: Grounding Performance
The t-SNE plot (Fig. 4) of Test Action embeddings for the base model shows clusters with LLM-suggested labels indicating semantic overlap (e.g., ’Navigation Control’ across multiple clusters). The grounding failure heatmap reveals that errors are distributed across the embedding space, not perfectly aligning with cluster boundaries. Notable high-failure regions overlap with clusters tentatively labeled Climate Control, Audio and Display Settings and User Interface Control, Device Settings. A distinct cluster associated with Navigation Control, User Interface Control resides in a predominantly low-failure region, suggesting these utterance types are generally grounded successfully.
@@
Using the same text embeddings, the fine-tuned model’s performance is visualized through its distinct failure heatmap (Fig. 5). When comparing Fig. 5 to Fig. 4, a noticeable reduction in both the extent and intensity of high-failure regions across the embedding space can be observed, supporting the results summarized in Table 3. Several clusters show significant improvement. For instance, the most distinct Cluster 3: Navigation Control, User Interface Control achieved a failure rate of close to zero. The dispersed Cluster 1: User Interface Control, Communication Management, Cluster 4: Climate Control, Vehicle Settings and Cluster 5: Climate Control, Vehicle Settings now predominantly reside in low-failure regions after fine-tuning, indicating that the model has learned to handle these automotive domain specific utterance types more reliably. However, some high-failure samples still persist at the boundaries of these clusters, suggesting that certain semantic regions or utterance formulations remain challenging even after fine-tuning.
@@
5.4.2 Expected Result Utterances: Verification and Grounding Performance
@@
The Expected Result utterances, representing a more complex task involving visual grounding and subsequent evaluation, also form clusters in the embedding space with themes like ’Settings’, ’Navigation’, and ’Display’. The base Molmo-7B failure heatmap (Fig. 6) for this task shows widespread and often intense high-failure regions, particularly concentrated in areas associated with User Interface Settings, Search Functionality and Audio Control, Radio Management.
@@
The failure heatmap (Fig. 7) for the fine-tuned model on the Expected Result tasks shows that high-failure regions persist across many parts of the embedding space. While a visual comparison with the base Molmo model might suggest a subtle reduction in the overall failure density, the improvement appears less pronounced than observed for the Test Action task.
@@
Fine-tuning provided only modest improvements for the more complex Expected Result verification and grounding task. The widespread persistence of high-failure regions suggests that accurately interpreting and verifying diverse state descriptions based on visual evidence remains a significant challenge.
5.4.3 Summary
The t-SNE visualization of a fixed text embedding space overlaid with task-specific VLM failure rates provided valuable insights by decoupling utterance semantics from model performance. The analysis indicated tangible benefits of fine-tuning for more direct tasks such as the Test Action grounding, while suggesting that more complex tasks like the Expected Result verification still present significant challenges. Additionally, the t-SNE plots highlighted variations in performance across different sub-domains. The model performed well in grounding UI elements commonly found in desktop, mobile, and web environments but showed reduced effectiveness in automotive-specific functionalities such as Advanced Driver-Assistance Systems (ADAS) or climate control, for instance, distinguishing between adjusting the driver’s and passenger’s temperatures. This underscores the general need for further domain-specific data and training strategies.

### v3

@@ 3.3 Model Fine-Tuning
For efficient training of baseline models, LoRa [12] was utilized. This parameter-efficient fine-tuning method injects small, trainable matrices into the model’s layers, allowing for significant improvements without retraining the entire large model. This approach offers benefits such as accelerated training, reduced training data needs, a smaller storage footprint for the adapters, lower memory requirements, and the potential for dynamic adapter swapping in deployment, which eliminates the need to load different large models for various tasks.
As previously established, Molmo-7B-D-0924 was chosen as the foundation for ELAM due to its strong performance on the AutomotiveUI-Bench-4k dataset among evaluated baseline models. This foundation provided a robust starting point. Fine-tuning with LoRA then enabled efficient adaptation to the specific tasks of automotive UI validation, leveraging a synthetically generated dataset that explicitly includes both Test Actions and Expected Results with their corresponding passed/failed states. This fine-tuning specifically aims to imbue ELAM-7B with the evaluation capabilities lacking in prior interaction-focused LAMs, and to shift the learned domain to automotive UI. Further details regarding this fine-tuning can be found in Section 5.1.
@@ 4 Open Source Infotainment Validation Dataset: AutomotiveUI-Bench-4K
Despite the availability of numerous datasets for benchmarking LAMs in desktop, mobile, and web UIs [29, 37, 36, 6, 43, 2, 8], a significant domain gap remains in their application to automotive UIs. To address this limitation, AutomotiveUI-Bench-4K is introduced and released as a new dataset specifically designed for this domain. This dataset comprises 998 images with 4,208 annotations from 15 brands, including Audi, BMW, Ford, Porsche, and Tesla. This diverse collection of modern vehicle UIs was annotated by experienced professional software test engineers specializing in Human-Machine Interface (HMI) evaluation. In addition to images from automotive OEMs, the dataset also incorporates screenshots of Apple CarPlay and Google Android Auto. Images from OEMs were captured in 4K resolution using cameras and sourced from either a research vehicle fleet or through collaboration with local car dealerships. To ensure consistent image perspective and focus on the relevant UI elements, the OEM images were rectified and cropped to the active display area. Android Auto and CarPlay screenshots were directly exported. The systems featured in this dataset are exclusively touchscreen-based. While some systems include auxiliary physical buttons, these were not considered in the analysis due to their primarily static functionality. This focused approach is justified by the increasing ubiquity of touchscreen interfaces as the dominant HMI in modern vehicles. By concentrating solely on this technology, the dataset provides a unified corpus for analysis. The resulting data homogeneity is ideal for robust comparative analysis and the development of VLMs specifically for this contemporary UI paradigm.
Although the dataset primarily comprises systems from recent model years, the system representing the earliest model year is a 2018 Volkswagen Tiguan. Languages that were included are German and English. Instructions and evaluation requests are written in English. German text was either translated or directly quoted. Real-world test engineering documentation often exhibits a pattern of short sentences and occasional grammatical deviations. Therefore, to create a dataset that realistically represents this domain, it is crucial to preserve these stylistic features in the preprocessing steps. Table 1 provides further information about the data distribution. The actions and results are unrelated, as test actions can lead to different UI states, and screen trajectories are not considered. There are two annotation classes in the dataset:
@@
    Test Action: Describes an interaction with a single UI element as an imperative sentence (”set A/C to max” in Figure 3).
@@
    Expected Result: Represents a testable expectation, defined per image, focusing on the required appearance, context, or state of UI elements within that image. This expectation is evaluated to determine a passed/failed outcome. (”Passenger’s climate zone is synced to driver” in Figure 3)
@@ 5 Experiments
This section details the experimental procedure for fine-tuning and evaluating ELAM. The following subsections cover the specific configuration used for training, the performance of the fine-tuned model against several baselines on the AutomotiveUI-Bench-4K and ScreenSpot datasets. The impact of different training data preprocessing strategies is explored through ablation studies, the relationship between language semantics and model performance is investigated using t-SNE, and a detailed visual error analysis is conducted.
5.1 Experimental Configuration
To train ELAM, LoRa was applied to all linear layers within the Molmo architecture. Training was conducted for 2 epochs with a global batch size of 128, achieved using a local batch size of 8 and gradient accumulation. A learning rate of $1\mathrm{e}{-4}$ was employed. An optimal LoRa rank $r=64$ was selected based on an ablation study (Table 5), with the $\alpha$ parameter set to $r$ and a fixed dropout rate of $0.05$. ModelScope’s SWIFT framework [42] was used to conduct fine-tuning.
Training data for ELAM was generated using the data pipeline described in Section 3.2. Gemini (gemini-2.0-flash-thinking-exp-01-21) was selected for training dataset generation due to its superior adherence to structured output compared to other models. The detailed prompts used for generating the training datasets are displayed in Sections A.3.1, A.3.2 and A.3.3.
Generated text for reasoning, test actions, and expectations was rephrased using a smaller model (gpt-4o-mini-2024-07-18). The synthetic dataset comprises 17,708 annotations across 6,230 images. This includes 5,952 Test Actions, 6,190 passed and 5,566 failed Expected Results. Prompt templates utilized for training, evaluation, and inference are detailed in Section A.2. Model training was performed on two NVIDIA H100 80 GB PCIe GPUs. For evaluation experiments on the AutomotiveUI-Bench-4K dataset, an NVIDIA GeForce RTX 4090 GPU with 24 GB VRAM was used.
Performance of the models was assessed using accuracy metrics, all expressed as percentages. For the visual grounding task, two distinct accuracy metrics were employed: $TA_{vg}$ represents the visual grounding accuracy for Test Actions, while $ER_{vg}$ denotes the visual grounding accuracy for Expected Results. For both $TA_{vg}$ and $ER_{vg}$, accuracy was determined by verifying that the generated point or the centroid of the predicted box was contained within the annotated ground truth box. Furthermore, for the evaluation of Expected Results, $ER_{evl}$ was used, representing the classification accuracy (passed or failed). To facilitate a fine-grained analysis of language-specific performance, the image dataset was categorized into English and German subsets. Corresponding language-specific metrics are indicated by the superscripts EN and DE for English and German UIs, respectively. In cases where a generated evaluation could not be parsed from the model’s response, it was assigned the inverse of the ground truth value.
@@ 5.2 Evaluating Performance with AutomotiveUI-Bench-4K
The performance of various baseline and fine-tuned models is presented in Table 2. ELAM demonstrates a notable improvement in localization as indicated by $TA_{vg}$ (+16.3%) and $ER_{vg}$ (+6.1%) compared to the baseline model. Specifically for evaluation, the fine-tuned model achieved higher accuracy $ER_{evl}$ (+11.3%), precision (+8.9%), and recall (+6.4%) (Figure 4). This enhancement is primarily driven by a significant reduction in both false positives and false negatives, indicating fine-tuning effectively improved the model’s ability to correctly identify passed tests while simultaneously reducing misclassifications of failed tests as passed, and vice versa.
The findings suggest that models within the 7B parameter range exhibit no significant performance degradation when processing images with predominantly German text. A significant challenge for most baseline models was the generation of consistently parseable output. For example, InternVL2.5 exhibited good evaluation performance. However, it rarely generated bounding box coordinates. TinyClick and UGround were evaluated exclusively on their localization performance, as they lack evaluative capabilities. To explore the possibility of faster inference through task distribution across models, TinyClick was selected for LoRa-fine-tuning, given its small size of only 0.27 billion parameters. This approach (LAM-270M), however, did not produce performance on par with Molmo-7B, as shown in Table 2.
A notable benefit of ELAM is its accessibility for deployment. Unlike many larger state-of-the-art VLMs, the 7 billion parameter ELAM can be deployed on a current consumer-grade NVIDIA GPU with at least 24 GB VRAM. The invoke time has been recorded as an average of 2.4 seconds for test actions and 3.4 seconds for expected results when running on an NVIDIA GeForce RTX 4090 GPU, using the default Hugging Face Transformers backend and full resolution images of AutomotiveUI-Bench-4K.
To assess ELAM’s potential on the AutomotiveUI-Bench-4K dataset, an evaluation was conducted by an experienced quality assurance engineer. This expert achieved improvements of +6.9% $TA_{vg}$, +8,9% $ER_{vg}$, and +15.0% $ER_{evl}$ compared to ELAM. This result provides a high-performance human expert benchmark and suggests that enhanced outcomes can be accomplished through additional training.
This analysis with a human expert was instrumental in identifying inherent limitations within the synthetic dataset itself. The expert’s higher performance serves as an important indicator for data-related issues. For instance, the boxes may be insufficiently sized for visual grounding, or screen elements may not be sufficiently described, highlighting potential areas for dataset improvement. A detailed analysis is conducted in Section 5.6.
Table 2: Evaluation on AutomotiveUI-Bench-4K. $TA_{vg}$ and $ER_{vg}$ represent visual grounding accuracies for Test Actions and Expected Results, respectively. $ER_{evl}$ denotes the classification accuracy (passed/failed) for Expected Results. Language specific metrics are indicated by the superscipts EN and DE for English and German UIs. All values are accuracies in percent.
@@
(a) Molmo-7B-D-0924 (baseline)
(b) ELAM-7B
Figure 4: Confusion matrices for the Expected Results evaluation classification in AutomotiveUI-Bench-4K, with values normalized to percentages.
@@ 5.3 Evaluating Generalizability with ScreenSpot
To determine if the fine-tuning resulted in overfitting to the automotive UI domain, ELAM-7B was applied to the ScreenSpot dataset [4] and compared to SeeClick [4], ShowUI [20], OS-Atlas-Base-7B [37], Molmo-7B-D-0924 /Molmo-72B-0924 [5], and Qwen2-VL-based UGround-V1-7B [10], as shown in Table 3. The fine-tuning procedure was found not to impair the generalizability of Molmo. The average score per category was improved by 5.6% compared to its baseline (Molmo-7B-D-0924) and by 2.2% compared to Molmo-72B-0924.
Table 3: $TA_{vg}^{(EN)}$ results on ScreenSpot for selected models taken from [11]
@@ 5.4 Ablation Study
Ablation studies were performed to evaluate the contributions of simple reasoning and rephrasing modules integrated into the synthetic data generation pipeline. Training runs were conducted using distinct preprocessing strategies, the configurations of which are summarized in Table 4. The prompts from Section A.2 were used for experiments that incorporate reasoning, whereas the phrase “Think step by step, conclude” was replaced by “Determine” for runs that did not utilize reasoning. A LoRa-rank of 64 was chosen to fine-tune all linear layers for the ablation experiments. It was demonstrated that the inclusion of reasoning significantly enhances performance, as evidenced by improvements in both visual grounding ($TA_{vg},ER_{vg}$) and evaluation ($ER_{evl}$) accuracy. The critical role of reasoning for robust generalization to real-world data in potential future deployments is highlighted by this observation.
To optimize the LoRa configuration, a separate ablation study focusing on the rank parameter $r$ was conducted, the results of which are presented in Table 5. The data was preprocessed utilizing reasoning and rephrasing. This analysis determined that a rank of 64 yielded optimal performance. In all LoRa-based experiments, the alpha parameter was set to $r$, and the LoRa-dropout rate was fixed at $0.05$.
@@
To investigate the relationship between the semantic representation of user utterances and VLM performance, particularly within automotive subdomains, two distinct task types were analyzed: Test Actions and Expected Results, as outlined in Section 4. For both types, text embeddings of the utterances were generated using the nomic-embed-text-v1.5***nomic-embed-text-v1.5 is an improved variant of the nomic-embed-text-v1[26] model. It utilizes Matryoshka Representation Learning for flexible dimensionality reduction with minimal loss while supporting up to 8,192 tokens. The model is optimized for search, clustering, and classification in production. model. Their structure was then visualized using t-SNE dimensionality reduction [31]. Unsupervised k-Means clustering ($k=8$) was applied to the high-dimensional embeddings to identify potential thematic groups. Cluster labels were generated post-hoc via a LLM (gpt-4o-2024-11-20) interpretation of sampled utterances. The t-SNE projections were enhanced with a heatmap that visualizes the local failure rate for each designated target for Test Actions and both evaluation and grounding for Expected Results. This failure rate was calculated per grid cell based on evaluation results. Red indicates high failure density, and white indicates low failure density. This approach facilitates the visualization of how VLM performance, for both the base Molmo-7B and ELAM-7B models, varies across the semantic landscape defined by the text embeddings.
@@ 5.5.1 Test Action Utterances: Grounding Performance
Clusters with LLM-suggested labels indicating semantic overlap (e.g., ’Navigation Control’ across multiple clusters) are observed in the t-SNE plot (Figure 5) of Test Action embeddings for the base model. The grounding failure heatmap indicates that errors are distributed across the embedding space, without perfectly aligning with cluster boundaries. Notable high-failure regions overlap with clusters tentatively labeled Climate Control, Audio and Display Settings and User Interface Control, Device Settings. A distinct cluster associated with Navigation Control, User Interface Control resides in a predominantly low-failure region, which suggests that these utterance types are generally grounded successfully.
@@
Utilizing the same text embeddings, the fine-tuned model’s performance is visualized through its distinct failure heatmap (Figure 6). When comparing Figure 6 to Figure 5, a noticeable reduction in both the extent and intensity of high-failure regions across the embedding space can be observed, supporting the results summarized in Table 2. Several clusters show significant improvement. For instance, the most distinct Cluster 3: Navigation Control, User Interface Control exhibited a failure rate close to zero. The dispersed Cluster 1: User Interface Control, Communication Management, Cluster 4: Climate Control, Vehicle Settings and Cluster 5: Climate Control, Vehicle Settings now predominantly reside in low-failure regions after fine-tuning, indicating improved reliability in handling these automotive domain-specific utterance types. However, some high-failure samples still persist at the boundaries of these clusters, suggesting that certain semantic regions or utterance formulations remain challenging even after fine-tuning.
@@
5.5.2 Expected Result Utterances: Evaluation and Grounding Performance
@@
The Expected Result utterances, representing a more complex task involving visual grounding and subsequent evaluation, are also observed to form clusters in the embedding space with themes such as “Settings”, “Navigation”, and “Display”. The Molmo-7B failure heatmap (Figure 7) for this task exhibits widespread and often intense high-failure regions, particularly concentrated in areas associated with User Interface Settings, Search Functionality and Audio Control, Radio Management.
@@
The failure heatmap (Figure 8) for the fine-tuned model on the Expected Result tasks indicates that high-failure regions persist across many parts of the embedding space. While a visual comparison with the base Molmo model may suggest a subtle reduction in the overall failure density, the improvement appears less pronounced than that observed for the Test Action task.
@@
Fine-tuning yielded only modest improvements for the more complex Expected Result evaluation and grounding task. The widespread persistence of high-failure regions indicates that accurately interpreting and verifying diverse state descriptions based on visual evidence remains a significant challenge.
5.6 Visual Error Analysis
5.6.1 Visual Error Analysis: Test Actions
A detailed analysis of the ELAM’s failures in the AutomotiveUI-Bench-4K evaluation was conducted to identify the areas where the model continues to demonstrate deficiencies. To begin with, the Test Actions were examined. A comprehensive manual review and classification of all error cases was conducted. Examples and explanations for all identified categories can be found in the appendix in Section A.4. Table 6 summarizes the distribution of error categories for Test Actions.
The following categories were identified:
  1. 1.
    The bounding boxes utilized for the purpose of evaluating the result were of insufficient size.
  2. 2.
    The element to be determined was not adequately described.
  3. 3.
    The subject of interest was a specialized automotive icon.
  4. 4.
    The provided description was adequate; however, the presence of two similar elements resulted in confusion.
  5. 5.
    The model proved incapable of establishing a connection between two elements.
  6. 6.
    The model demonstrated confusion between driver side and passenger side, left and right, or increase and decrease.
  7. 7.
    The model was incapable of counting.
  8. 8.
    Not assignable to any category.
Table 6: Analysis of error categories based on Test Actions in AutomotiveUI-Bench-4K.
Category # Category Description Error Share in %
1. Box too small 17.0
2. Ambiguous or poor description of the test action 23.7
3. Special icons from the vehicle domain 19.7
4. Clear description, but very similar elements confused the model 11.5
5. Connection between elements could not be established 6.8
6. Driver/passenger, left/right, up/down distinctions 4.8
7. Elements must be counted 3.1
8. Miscellaneous 13.4
The findings underscore the importance of ensuring clear and precise phrasing in the test utterances used by developers and testers, as this practice can mitigate a range of potential issues. Analysis also showed that the system’s recognition of vehicle-specific icons, in particular, remains suboptimal, suggesting that performance would benefit from additional training data. A more fundamental challenge identified is the system’s difficulty in differentiating between highly similar elements and establishing connections between them. The creation of specialized training datasets to address these specific scenarios represents a promising avenue for future research.
5.6.2 Visual Error Analysis: Expected Results
An analysis of the ELAM’s failures of grounding and evaluating Expected Results in the AutomotiveUI-Bench-4K evaluation was conducted in the same manner as the examination of the Test Actions. The errors were manually reviewed and categorized. The identified categories were similar to those found in the examination of the Test Actions, though some differences were noted. Table 7 summarizes the distribution of error categories for Expected Results. Examples and explanations for all identified categories can be found in the appendix in Section A.5.
The following categories were identified:
  1. 1.
    The bounding boxes utilized for the purpose of evaluating the result were of insufficient size.
  2. 2.
    Multiple areas display the expected result.
  3. 3.
    The model is asked to check if options of a certain menu are visible. However the box was not drawn around the menu option, but around the menu heading instead. This could be considered as a special case of category 1 or 2.
  4. 4.
    The subject of interest was a specialized automotive icon.
  5. 5.
    The provided description was adequate, however the presence of two similar elements resulted in confusion.
  6. 6.
    Not assignable to any category.
Table 7: Analysis of error categories based on Expected Results in AutomotiveUI-Bench-4K.
Category # Category Description Error Share in %
1. Box too small 26.4
2. Multiple areas display the expected result 15.2
3. Box not around option, when asked if options are shown 15.2
4. Special icons from the vehicle domain 8.8
5. Clear description, but very similar elements confused the model 4.8
6. Miscellaneous 29.6